OpenStudy (cutiecomittee123):

What is the best approximation of the projection of (2, 6) onto (5, -1)

OpenStudy (mathmale):

Is this vector calculus? If so, please re-write your two numerical expressions as vectors. Otherwise other people will think you're discussing points (x,y).

OpenStudy (cutiecomittee123):

This is what is given in the question, I am not sure how to write them otherwise

OpenStudy (dumbcow):

try using notation <2,6> for vectors

OpenStudy (dumbcow):

reference: https://en.wikipedia.org/wiki/Vector_projection#Vector_projection

OpenStudy (dumbcow):

first determine magnitude of both vectors \[|<2,6>| = \sqrt{2^2 + 6^2} = 2 \sqrt{10}\] \[|<5,-1>| = \sqrt{5^2 + 1^2} = \sqrt{26}\] Next find dot product: \[\rightarrow (2*5) + (6*-1) = 4\] Determine unit vector of 2nd vector \[\rightarrow <\frac{5}{\sqrt{26}}, -\frac{1}{\sqrt{26}}>\] Now plug into projection equation: \[(2 \sqrt{10})(\frac{4}{2 \sqrt{10} \sqrt{26}}) <\frac{5}{\sqrt{26}},-\frac{1}{\sqrt{26}}>\] Finally the projection vector is: \[<\frac{10}{13}, -\frac{2}{13}>\]

OpenStudy (tkhunny):

"best approximation of the projection " What does that mean? It IS the projection. It's not an approximation.